OpenStudy (anonymous):

Solve: –(3x – 2) = –4(x + 1) – 4 ?

Parth (parthkohli):

Do you know the distributive property? :)

Parth (parthkohli):

\(\Rightarrow -3x + 2 = -4x - 4 - 4\) Can you solve now?

OpenStudy (anonymous):

Yes! -10? thanks so much - what exactly did you do to get it into that form?

OpenStudy (anonymous):

He distributed

OpenStudy (anonymous):

-3x+4x=-4-4-2 => x= -10